OpenStudy (caitlinnr14):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers. please solve

OpenStudy (caitlinnr14):

show steps

OpenStudy (caitlinnr14):

anyone?? help

OpenStudy (anonymous):

arnt "real numbers" all numbers?

OpenStudy (caitlinnr14):

yeas

OpenStudy (caitlinnr14):

help please

OpenStudy (anonymous):

ummm.... 2x-4=12 i think

OpenStudy (anonymous):

The equation : \[\Large 2x+6=2(x+3)\] has all the real number as solution